OpenStudy (anonymous):

Complex numbers help! if ( x + iy )² = 4i find all the real values of x and y

hartnn (hartnn):

so, expand (x+iy)^2 can you ?

OpenStudy (anonymous):

\[x ^{2} - y ^{2} + 2xyi = 4 i\]

hartnn (hartnn):

correct

hartnn (hartnn):

now whats the real part of that ?

OpenStudy (anonymous):

x² - y²

hartnn (hartnn):

yes, and real part of right side =0 right ? so equating the real parts, we get x^2-y^2 =0 got this ?

OpenStudy (anonymous):

ohhhh x² - y² = 0

hartnn (hartnn):

yes

hartnn (hartnn):

now equate the imaginary part too

OpenStudy (anonymous):

2xyi = 4i

hartnn (hartnn):

so 2xy = 4 right ?

OpenStudy (anonymous):

yea but there is only two variables i've been struggling with this for a half an hour

hartnn (hartnn):

we have 2 equations and 2 variables so we can find them easily

hartnn (hartnn):

x^2-y^2 = 0 what do you get from this ?

OpenStudy (anonymous):